the length of a rectangle is twice as long as the width and its perimeter is 54 ft. find the dimensions of the rectangle. can you show me the steps.
Answer by oberobic(2304)

L = 2W
.
P = 2(L+W)
P = 54
.
substitute
.
2(L+W) = 54
2(2W + W) = 54
4W +2W = 54
6W = 54
W = 9
.
L = 2W
L = 2(9) = 18
.
Check the perimeter
.
2(18+9) = 2(27) = 54
Correct.
.
Answer: Length = 18; Width = 9.
.
Done.